OpenStudy (anonymous):

Verify the following identity: cot(x - pi/2)= -tanx

OpenStudy (anonymous):

i have this : tan (–x) = –tan x cot (pi/2-x)= tan plug in -x for x. = (x+pi/2)

OpenStudy (anonymous):

but im doing something wrong

OpenStudy (shamim):

\[\cot (x- \frac{ \pi }{ 2 })=\cot(-(\frac{ \pi }{ 2 }-x))=-\cot (\frac{ \pi }{ 2 }-x)=-tanx\]

OpenStudy (shamim):

actually\[\cot (\frac{ \pi }{ 2 }-x)=\tan x\]

OpenStudy (shamim):

because its 1st quadrant

OpenStudy (anonymous):

so i do -(pi/2-x)?

OpenStudy (shamim):

u should tell\[\cot(-(\frac{ \pi }{ 2 }-x))\]

OpenStudy (shamim):

\[=-\cot(\frac{ \pi }{ 2 }-x)\]

OpenStudy (shamim):

\[=-\tan x\]

OpenStudy (anonymous):

ok thanks!

OpenStudy (shamim):

welcome